The price of oil is $30 per barrel and the price elasticity is constant and equal to -0.5. An oil embargo reduces the quantity available by 20 percent. Use the arc elasticity formula to calculate the percentage increase in the price of oil.
1
Expert's answer
2015-07-25T00:00:42-0400
If he price of oil P = $30 per barrel and the price elasticity is constant and equal to E = -0.5, and if an oil embargo reduces the quantity available by 20 percent, the new price of oil using the arc elasticity formula will be P2 = 30*(1 + 0.2/|-0.5|) = 30*1.4 = $42.